A father is now 7 times as old as his son. In 5 years he will be 4 times as old as his son will be then. What are their present ages?
Found 2 solutions by maribeth_abella, ghall:
Answer by maribeth_abella(11)

Let x = son's present age
7x = father's age

E: 7x + 5 = 4(x + 5)
7x + 5 = 4x + 20
7x -4x = 20 - 5
3x = 15
x = 15/3
x = 5
A: x = son's age
7x = 7(5)= 35 - father's age
Checking: 7(5) + 5 = 4(5 + 5)
35 + 5 = 20 + 20
40 = 40

Answer by ghall(1)